# Knowledge Enhanced Conditional Imputation for Healthcare Time-series

### v3

@@ Abstract
We present an end-to-end architecture for managing complex missingness in multivariate time series derived from hospital electronic health records (EHRs). Our Conditional Self-Attention Imputation (CSAI) is a recurrent neural network architecture equipped with a number of techniques aiming to improve imputation accuracy by aligning the model with the subtle temporal and spatial dependencies typical of clinical data. CSAI a) utilises an attention-based hidden state initialisation to capture long- and short-range correlations within the time-series, b) incorporates a knowledge embedding technique to capture clinical data recording patterns and c) employs a non-uniform masking strategy to adapt its weights to the data’s temporal and cross-sectional missingness patterns. Extensive evaluation of three EHR benchmark data sets demonstrates that CSAI enhances the current state of the art efficacy in data restoration in addition to performance on downstream tasks. Furthermore, CSAI is integrated within the PyPOTS Python library for benchmarking, offering open and standardised benchmarking capabilities and ease of use for researchers.
keywords:
neural network imputation, healthcare time-series, neural architectures, multivariate time-series imputation, complex missingness
1 Introduction
@@
Multivariate time series obtained from electronic health records (EHRs) play a crucial role in predictive healthcare analytics [1], with many models developed to derive patient-specific insights from EHR data [2, 3]. However, EHR time series are characterised by their complexity, and designing successful downstream predictive models requires addressing the abundance of missingness in correlated variables recorded over time. The nature and timing of the data recorded in EHRs are governed by clinical and administrative decisions, leading to variations in the frequency of documenting vital signs, laboratory tests, and other healthcare measures. For example, heart rate is routinely monitored, while data on white blood cell counts are often scarce, as they are ordered only in specific clinical scenarios, such as suspected infection. In addition, the characteristics typical of EHR data underlie the abundance of correlations between the values and missingness patterns of features over time. For example, because hypertension is a known cause of kidney disease, not only are high blood pressure measurements expected along with abnormal creatinine levels, but EHR datasets generally show high correlations between the recording patterns of blood pressure and urine creatinine levels [4]. The inherent irregularities render more than 50% of EHR data missing not at random [5], with significant variations in missingness patterns between tasks and datasets [5]. These challenges create a complex landscape for imputation algorithms [6, 7, 8, 9] and highlight the importance of domain knowledge in modelling accuracy.
@@
Recent advances in deep learning have shown significant potential in addressing the complexities of EHR missingness. In particular, models based on Recurrent Neural Networks (RNNs), such as GRU-D [10], BRITS [11], and M-RNN [12], have proven effective in capturing EHR temporal and cross-sectional dependencies. Among these, BRITS has shown superior performance in datasets of varying characteristics [13, 14]. BRITS uses bi-directional RNN dynamics and treats missing data as variables of the bidirectional RNN graph, subsequently involving them in the backpropagation process. Using this strategy, missing values accumulate gradients in both forward and backward directions with consistency constraints, increasing the accuracy of estimated values.
@@
Despite advances, none of the existing models, including BRITS, account for domain-imposed disparities in feature-recording patterns, which can have critical implications on imputation accuracy. We show that incorporating a domain-informed strategy to capture feature correlations is particularly suited for our domain. In addition, the temporal correlations in the EHR data can span both short- and long-term intervals. For example, while monitoring diabetic patients, both short-term glucose fluctuations and long-term HbA1c levels must be accurately modelled, while cardiovascular risk assessment requires careful cross-sectional examination of biomarkers like cholesterol, glucose, and blood pressure. Although RNNs excel at capturing short-term temporal dependencies in a time series, they are less capable of capturing long-range correlations over time [7]. Finally, all current models rely on random masking as a pre-processing step to generate ground truths for evaluation purposes. This involves designating a randomly selected subset of complete values as missing, primarily creating missing completely at random (MCAR) scenarios and drastically oversimplifying the temporal and cross-sectional dependencies in real EHR time series.
@@
    Improved initialisation of hidden states to more effectively model temporal dynamics between data points. Our conditional hidden state initialisation leverages self-attention to account for long-term dynamics, complementing those already captured by the RNN and propagating them through the hidden states, thereby improving overall performance.
@@
    We present a domain-informed temporal decay function that determines the likelihood of missing values being linked to past observations based on clinical recording patterns in the dataset. The decay of a feature is used to adjust its associated attention mechanism, enabling a more fine-grained, feature-specific representation of temporal relationships within the data.
@@
    CSAI is integrated with a non-uniform masking strategy to selectively mask data in a way that reflects the inherent structured patterns of interdependencies within the dataset across time and features.
The missing data patterns in the EHR reflect the underlying data structures and convey clinically relevant information about the variables used in the analysis tasks. Aligning the imputation process with these characteristics can significantly impact the quality of downstream predictive models. CSAI’s contributions mark an important step toward enhancing imputation algorithms to better capture domain-specific patterns in EHR data, leading to improved performance in tasks like patient prognosis and clinical decision-making.
2 Related Work
The ability of RNNs to capture temporal evolution has made them an attractive choice for time-series imputation. The GRU-D model [10] introduced the use of temporal decay to recurrent imputation models, whereby the influence of past recordings on missing data decreases over time, with features imputed closer to default values if their last observation occurred a long time ago. Using this concept, GRU-D estimates missing values through a combination of the last recorded value and the global mean within a recurrent architecture. Although this approach alone has shown impressive performance in certain healthcare applications, its effectiveness has not been generalisable in other datasets [11], particularly in accounting for discrepancies in vital signs recording frequencies and laboratory tests.
Several extensions of GRU-D exist. M-RNN [12] and BRITS [11] use bi-directional RNNs to capture temporal dynamics in both forward and backward directions, improving imputation accuracy. Both M-RNN and BRITS combine RNNs with a feed-forward component to capture additional correlations across features. However, M-RNN treats imputed values as fixed constants while BRITS treats missing values as variables, updating them through backpropagation and demonstrating superior performance across various domains [13, 15].
@@ 2 Related Work
3 Terminology and Background
3.1 Incomplete Multivariate Time-series Representation
@@
3.2 Overview of the BRITS Backbone
BRITS exploits cross-sectional and temporal correlations in multivariate time series through two dedicated components, a fully connected regression module and a recurrent module, respectively. It also relies on the notion of temporal decay (Eq.(4)) to capture temporal correlations and adjust their influence on missing values based on temporal distance. The BRITS architecture accepts a time series $\boldsymbol{X}$, such that missing values within an observation $\boldsymbol{x_{t}}\in\boldsymbol{X}$ are managed via a corresponding masking vector $\boldsymbol{m_{t}}$ and a time gap vector $\boldsymbol{\delta_{t}}$ as shown at the bottom of Fig.2.
Because an observation $\boldsymbol{x_{t}}$ may contain missing values, it cannot be directly ingested by the two BRITS components. Instead, BRITS uses the notion of decay to derive estimates for missing cells. Temporal decay dictates that the strengths of the correlations are inversely related to the time-gap $\boldsymbol{\delta_{t}}$ and is formalised by the decay factor $\gamma_{t}\in(0,1]$ (Eq.(4)). In the equation, $\boldsymbol{W}_{\gamma}$ and $\boldsymbol{b}_{\gamma}$ are parameters. The calculated decay factor $\gamma_{t}$ for the hidden state is then used to transform $\boldsymbol{h}_{t-1}$ to the decayed hidden state, $\boldsymbol{\hat{h}}_{t-1}$ as in Eq.(5). BRITS uses the decayed hidden state to find the historical estimation $\boldsymbol{\hat{x}_{t}}$, of $\boldsymbol{x_{t}}$ as in Eq.(6). Using masking vector $\boldsymbol{m_{t}}$ obtained from $\boldsymbol{M}$, $\boldsymbol{\hat{x}_{t}}$ is then used to replace the missing values of $\boldsymbol{x_{t}}$, yielding the complement vector $\boldsymbol{x_{t}^{h}}$, which embeds temporal missingness patterns to be fed into subsequent BRITS components (Eq.(7)).
@@ 3.2 Overview of the BRITS Backbone
$\boldsymbol{\hat{x}_{t}}$ $=\boldsymbol{W_{x}}\boldsymbol{\hat{h}_{t-1}}+\boldsymbol{b_{x}}$ (6)
@@
Figure 2: The BRITS backbone processes
BRITS explores cross-sectional correlations within a single observation through a fully connected layer, generating $\boldsymbol{x_{t}^{f}}$, a feature-wise approximation of missing values (Eq.(8)). The concept of decay extends to the feature space, resulting in a learnable factor $\hat{\beta}{{}_{t}}$, taking into account the temporal decay and the masking vector (Eq.(9)-(10)). This integration produces the imputed matrix $\boldsymbol{x_{t}^{c}}$, effectively combining observed and imputed data (Eq.(11)-(12)).
@@
The final step of BRITS (Eq.(13)) updates the hidden state via its RNN component, leveraging various indicators to learn functions of past observations. Bidirectional recurrent dynamics is used to address slow convergence by using information in the backward direction.
The entire process is shown in Fig.2. BRITS bidrectional architecture combined with its decay mechanism represents a significant advancement for multivariate time series imputation, addressing the challenge of temporal correlations at a level of sophistication unmatched by traditional methods.
4 Methodology
In this section, we describe the modifications proposed to the BRITS architecture to incorporate a) a domain-informed temporal decay functionality, and b) the ability to capture long-range correlations via a transformer-based conditional hidden state initialisation. In addition, we describe a novel non-uniform masking strategy to explicitly model non-random missingness.
4.1 Domain-informed Temporal Decay
The BRITS decay function Eq.(4) is strictly dependent on temporal proximity, dynamically adjusting the contribution of a past observation to a missing value based on the length of the time gap between the two. Although this decay mechanism effectively captures the intuition that more recent observations carry greater diagnostic value, it overlooks the discrepancies in the clinical recording patterns of different features. This can be illustrated in the example below.
@@ 4.1 Domain-informed Temporal Decay
Let features $f_{1}$ and $f_{2}$ be two features representing heart rate (HR) and systolic blood pressure (SBP) respectively; HR is typically recorded more frequently than SBP. BRITS is presented with an observation $\boldsymbol{x_{t}}$ with missing HR and SBP values. The time gap vector shows $\boldsymbol{\delta^{f_{1}}_{t}}=2$ and $\boldsymbol{\delta^{f_{2}}_{t}}=7$. According to the principle of temporal decay, BRITS will ignore the typical recording patterns of the two features and the last observation for $f_{2}$ will incorrectly have less weight on the imputed value than that of $f_{1}$.
Our new decay mechanism prioritises recent observations while accounting for the natural variability in healthcare data collection. In addition to using the observed time gap $\boldsymbol{\delta^{d}_{t}}$, we modify the decay function to incorporate the expected time gap $\boldsymbol{\tau}$ between two recordings of a given feature. $\boldsymbol{\tau}$ for a feature is the overall median of all gap lengths between any two consecutive feature recordings in the dataset. The adjusted decay function only reduces the weight assigned to an observation if its last observed value lies outside $\boldsymbol{\tau}$ and allocates more weight to observations closer to the expected time gap and less to those further away, as given below:
$A_{t}=\exp\left(-\max(0,\boldsymbol{W_{\gamma}}\boldsymbol{(\delta_{t}-\tau)}+\boldsymbol{b_{\gamma}})\right)$ (14)
This ensures that the attention peaks when the time gap $\boldsymbol{\delta_{t}}$ closely matches $\boldsymbol{\tau}$, and declines as the difference between $\boldsymbol{\delta_{t}}$ and $\boldsymbol{\tau}$ increases.
@@
For the same features $f_{1}$ and $f_{2}$, examining the dataset shows that the median time gaps for the two features are $\boldsymbol{\tau_{1}}=2$ for HR, compared to $\boldsymbol{\tau_{2}}=10$ for SBP respectively. These medians reflect the fact that HR is much more frequently monitored than SBP in routine clinical practice. Here, because the past recordings fall within the respective time gaps for both features, the past values will be assigned equal weights to the missing values. In this way, the less frequently measured SBP (with a median gap of 10) is not unfairly penalised.
4.2 Hidden State Initialisation
The hidden state of the recurrent component of BRITS are not generated through the raw input. Instead, they receive incomplete data with missingness indicators for imputation, i.e. $\boldsymbol{\hat{h}}$ replaces $\boldsymbol{h}$ in Eq.(5). However, the hidden states in BRITS are initialised with zeros. Since information decay starts at the first step, zeros $\boldsymbol{h_{0}}$ will cause $\boldsymbol{x^{h}_{1}}$ and $\boldsymbol{x^{f}_{1}}$ fully rely on neural network parameters, then accumulate.
Instead, our approach makes use of the last observed data point and a decay attention mechanism to generate a conditional hidden state distribution $\boldsymbol{q}(\boldsymbol{h}_{\text{init}}|\boldsymbol{x}_{\text{last\_obs}},\boldsymbol{A}(\boldsymbol{\delta_{t}},\boldsymbol{\tau}))$ within the model distribution $\boldsymbol{p_{\theta}}(\boldsymbol{x_{t}})$. This strategy is designed to provide a more contextually rich starting point for the model, thus enhancing the effectiveness of subsequent imputations. In addition, unlike in BRITS where the decay factor is applied directly to the previous hidden state Eq.(5) and Eq.(9), our approach uses the decay factor to modulate the attention mechanism. This modification allows for a more fine-grained and feature-specific generation of temporal relationships within the data, enhancing the model ability to adapt to varying temporal dynamics across different features. Our approach is described in detail below.
[…]
Input projection and positional encoding At each time step $\boldsymbol{s_{t}}$, the last observation $\boldsymbol{x_{\text{last\_obs}}}\in\mathbb{R}^{T\times D_{\text{feature}}}$ from input time-series data $\boldsymbol{x_{t}}$ undergo transformation alone with $\boldsymbol{A}(\boldsymbol{\delta_{t}},\boldsymbol{\tau})\in\mathbb{R}^{T\times D_{\text{feature}}}$ from Eq.(14). The Positional Encoding module and an Input Projection module are used to transform them into $\mathbb{R}^{T\times D_{\text{model}}}$, defined as:
@@ 4.2 Hidden State Initialisation
$\boldsymbol{A^{{}^{\prime}}}(\boldsymbol{\delta_{t}},\tau)$ $=\text{PosEncoder}(\text{InputProj}(\boldsymbol{A}(\boldsymbol{\delta_{t}},\boldsymbol{\tau}))$ (16)
[…]
Transformer enhancement The concatenated input for the Transformer Encoder $\boldsymbol{C_{\text{in}}}$ is formed by combining transformed last observation and decay attention and processed through multiheaded self-attention (MSA), layernorm (LN), and feed-forward networks (FFN).
$\boldsymbol{C}_{\text{in}}$ $=\text{Concat}(\boldsymbol{x^{{}^{\prime}}_{\text{last\_obs}}},\boldsymbol{A^{{}^{\prime}}}(\boldsymbol{\delta_{t}},\boldsymbol{\tau}))$ (17)
@@
[…]
Information scaling to initialise hidden state The enhanced output of the transformer Encoder is scaled to the dimensions required by the hidden state $\boldsymbol{h}_{init}$ through alternating 1D convolution.
@@
where Eq.(19) transform $\boldsymbol{C}_{\text{out}}$ from $\mathbb{R}^{2L\times d_{\text{model}}}$ to $\mathbb{R}^{2L\times d_{\text{hidden}}}$ and produces $\boldsymbol{H_{1}}$, Eq.(20) further scale $\boldsymbol{H_{1}}$ to generate the initialized Hidden State $\boldsymbol{h}_{\text{init}}$. The CSAI architecture is shown in Fig.3.
Figure 3: The CSAI architecture, which equips BRITS with self-attention. CSAI begins with an input embedding layer, transforming raw data into a dense representation. Positional embedding is applied next, adding temporal context to the input, enabling the model to capture time dependencies. These embeddings are concatenated to form a unified input that integrates the data’s temporal structure. The input is then processed through multi-head attention, normalisation, and feed-forward layers, with decay attention mechanisms modulating the importance of previous observations. The final output is used to initialise hidden states for subsequent recurrent layers, accounting for both temporal dynamics and domain-specific variability in recording patterns.
4.3 Non-Uniform Masking Strategy
Our non-uniform masking strategy fundamentally diverges from traditional approaches by leveraging the inherent missingness distribution. This strategy is predicated on the principle outlined in the introduction, whereby the probability of missingness in healthcare data is not uniformly distributed across features. Instead, it varies according to healthcare practices and patient conditions. By incorporating this variability into our masking process, we aim to create a more realistic and representative model of the missing data patterns.
The core of our algorithm involves generating non-uniform masking probabilities guided by the missingness distributions of each feature. Another essential aspect is the manual adjustment factor, designed to fine-tune the masking process. This factor is adaptively set based on the observation frequency per feature, allowing for a tailored masking strategy. This adaptive approach aims to achieve a balance between avoiding overfitting in sparse data scenarios and ensuring sufficient learning signals in data-rich contexts. It is designed to optimise the model’s ability to discern underlying patterns and relationships, enhancing its predictive accuracy and generalisation across diverse healthcare data patterns. For a given feature $d$, the non-uniform masking probability $\boldsymbol{P_{\text{nu}}}(d)$ is determined as follows:

### v4

@@ Abstract
We introduce the Conditional Self-Attention Imputation (CSAI), a novel recurrent neural network architecture designed to address the challenges of complex missing data patterns in multivariate time series derived from hospital electronic health records (EHRs). CSAI extends the current state-of-the-art neural network-based imputation methods by introducing key modifications specifically adapted to EHR data characteristics, namely: a) an attention-based hidden state initialisation technique to capture both long- and short-range temporal dependencies prevalent in EHRs, b) a domain-informed temporal decay mechanism to adjust the imputation process to clinical data recording patterns, and c) a non-uniform masking strategy that models non-random missingness by calibrating weights according to both temporal and cross-sectional data characteristics. Comprehensive evaluation across four EHR benchmark datasets demonstrate CSAI’s effectiveness compared to state-of-the-art neural architectures in data restoration and downstream predictive tasks. Additionally, CSAI is integrated within PyPOTS, an open-source Python toolbox designed for machine learning tasks on partially observed time series. This work significantly advances the state of neural network imputation applied to EHRs by more closely aligning algorithmic imputation with clinical realities.
I Introduction
@@
Multivariate time series obtained from electronic health records (EHR) play a crucial role in predictive healthcare analytics to derive patient-specific insight [1, 2, 3]. However, EHR time series are characterised by their complexity, and designing successful downstream predictive models requires addressing the abundance of missingness within correlated variables recorded over time. EHR missingness is a direct consequence of clinical and administrative workflows, which inherently lead to variations in the frequency and timing of documenting vital signs, laboratory tests, and other healthcare measures. For example, heart rate is routinely monitored, while data on white blood cell counts are ordered in specific clinical scenarios, such as suspected infection. Furthermore, the structure of EHR data gives rise to numerous correlations between the feature values and their missingness patterns over time. For example, because hypertension is a known cause of kidney disease, not only are high blood pressure measurements expected along with abnormal creatinine levels, but EHR datasets generally show high correlations between the recording patterns of blood pressure and urine creatinine levels [4]. These inherent irregularities render more than 50% of EHR data missing not at random [5], with significant variations in missingness patterns between tasks and datasets [5]. These challenges create a complex landscape for imputation algorithms [6, 7, 8, 9, 10, 11, 12] and highlight the importance of domain knowledge in improving modelling accuracy.
@@
Recent advances in deep learning have shown significant potential in addressing the complexities of EHR missingness. In particular, models based on recurrent neural networks (RNNs), such as GRU-D [13], BRITS [14], and M-RNN [15], have proven effective in capturing the temporal and cross-sectional dependencies embedded within EHR data. Among these, BRITS has shown superior performance in datasets of varying characteristics [16, 17]. BRITS uses bi-directional RNN dynamics and treats missing data as variables of the bidirectional RNN graph, subsequently involving them in the backpropagation process.
@@
Despite advances, none of the existing models, including BRITS, account for domain-imposed disparities in feature-recording patterns, which can have critical implications on imputation accuracy. We show that incorporating a domain-informed strategy to capture data recording patterns is particularly suited for our domain. In addition, the temporal correlations in the EHR data can span both short- and long-term intervals. For example, while monitoring diabetic patients, both short-term glucose fluctuations and long-term HbA1c levels must be accurately modelled, while cardiovascular risk assessment requires careful cross-sectional examination of biomarkers like cholesterol, glucose, and blood pressure. Although RNNs excel at capturing short-term temporal dependencies in a time series, they are less capable of capturing long-range correlations over time [8]. Finally, all current models rely on random masking as a pre-processing step to generate ground truths for evaluation purposes. This involves designating a randomly selected subset of complete values as missing, primarily creating missing completely at random (MCAR) scenarios and drastically oversimplifying the temporal and cross-sectional dependencies in real EHR time series.
@@
    Improved initialisation of hidden states to more effectively model temporal dynamics between data points. Our conditional hidden-state initialisation leverages self-attention to account for long-term dynamics, complementing the short-term dependencies already captured by the RNN and propagating them through the hidden states, thereby improving overall performance.
@@
    We present a domain-informed temporal decay function that determines the likelihood of missing values being linked to past observations based on clinical recording patterns derived from the dataset. The decay factor of a feature is used to adjust its associated attention mechanism, enabling a more fine-grained, feature-specific representation of temporal relationships within the data.
@@
    CSAI is integrated with a non-uniform masking strategy to selectively mask data in a way that reflects the naturally structured patterns of interdependencies within the dataset across time and features.
The missing data patterns in the EHR convey clinically relevant information about the variables used in subsequent tasks. Aligning the imputation process with these characteristics can significantly impact the quality of downstream predictive models. CSAI’s contributions better capture domain-specific patterns in EHR data during imputation, leading to improved performance in tasks such as patient-specific prognosis and decision-making.
II Related Work
The ability of RNNs to capture temporal evolution has made them an attractive choice for time-series imputation. The GRU-D model [13] introduced the use of temporal decay to recurrent imputation models, whereby the influence of past recordings on missing data decreases over time, with features imputed closer to default values if their last observation occurred a long time ago. Using this concept, GRU-D estimates missing values through a combination of the last recorded value and the global mean within a recurrent architecture. Although this approach alone has shown impressive performance in certain healthcare datasets, its effectiveness has not generalised in others [14], particularly with respect to accounting for discrepancies in vital signs recording frequencies and laboratory tests.
Several extensions of GRU-D exist. M-RNN [15] and BRITS [14] use bi-directional RNNs to capture temporal dynamics in both forward and backward directions, improving imputation accuracy. Both M-RNN afnd BRITS combine RNNs with a feed-forward component to capture additional correlations across features. However, M-RNN treats imputed values as fixed constants while BRITS treats missing values as variables, updating them through backpropagation. Using this strategy, missing values accumulate gradients in both forward and backward directions with consistency constraints, increasing the accuracy of estimated values. BRITS has demonstrated superior performance across various domains [16, 18].
@@ II Related Work
III Terminology and Background
III-A Incomplete Multivariate Time-series Representation
@@
III-B Overview of the BRITS Backbone
BRITS exploits cross-sectional and temporal correlations in multivariate time series through two dedicated components, a fully connected regression module and a recurrent module, respectively. The BRITS architecture accepts a time series $\boldsymbol{X}$, such that missing values within an observation $\boldsymbol{x_{t}}\in\boldsymbol{X}$ are managed via a corresponding masking vector $\boldsymbol{m_{t}}$ and a time gap vector $\boldsymbol{\delta_{t}}$ as shown at the bottom of Fig.2.
Because an observation $\boldsymbol{x_{t}}$ may contain missing values, it cannot be directly ingested by the two BRITS components. Instead, BRITS uses the notion of decay to derive estimates for missing cells. Temporal decay dictates that the strengths of the correlations are inversely related to the time-gap $\boldsymbol{\delta_{t}}$ and is formalised by the decay factor $\gamma_{t}\in(0,1]$, Eq.(4). In the equation, $\boldsymbol{W}_{\gamma}$ and $\boldsymbol{b}_{\gamma}$ are parameters. The calculated decay factor $\gamma_{t}$ for the hidden state is then used to transform $\boldsymbol{h}_{t-1}$ to the decayed hidden state, $\boldsymbol{\hat{h}}_{t-1}$ as in Eq.(5). BRITS uses the decayed hidden state to find the historical estimation $\boldsymbol{\hat{x}_{t}}$, of $\boldsymbol{x_{t}}$ as in Eq.(6). Using masking vector $\boldsymbol{m_{t}}$ obtained from $\boldsymbol{M}$, $\boldsymbol{\hat{x}_{t}}$ is then used to replace the missing values of $\boldsymbol{x_{t}}$, yielding the complement vector $\boldsymbol{x_{t}^{h}}$, which embeds temporal missingness patterns to be fed into subsequent BRITS components, Eq.(7).
@@ III-B Overview of the BRITS Backbone
$z\boldsymbol{\hat{x}_{t}}$ $=\boldsymbol{W_{x}}\boldsymbol{\hat{h}_{t-1}}+\boldsymbol{b_{x}}$ (6)
@@
BRITS explores cross-sectional correlations within a single observation through a fully connected layer, generating $\boldsymbol{x_{t}^{f}}$, a feature-wise approximation of missing values, Eq.(8). The concept of decay extends to the feature space, resulting in a learnable factor $\hat{\beta}{{}_{t}}$, taking into account the temporal decay and the masking vector, Eq.(9)-(10). This integration produces the imputed matrix $\boldsymbol{x_{t}^{c}}$, effectively combining observed and imputed data (Eq.(11)-(12)).
@@
The final step of BRITS updates the hidden state via its RNN component, leveraging various indicators to learn functions of past observations, Eq.(13). Bidirectional recurrent dynamics is used to address slow convergence by using information in the backward direction. The entire process is shown within the BRITS bidrectional architecture in Fig.2
Figure 2: The BRITS backbone process
IV Methodology
In this section, we describe the modifications proposed to the BRITS architecture to incorporate a) a domain-informed temporal decay functionality, b) the ability to capture long-range correlations via a transformer-based conditional hidden state initialisation, and c) a novel non-uniform masking strategy to explicitly model non-random EHR missingness. The section concludes with the resulting learning framework of CSAI.
IV-A EHR-Tailored BRITS Adaptations
Domain-informed Temporal Decay: The BRITS decay function Eq.(4) is strictly dependent on temporal proximity, dynamically adjusting the contribution of a past observation to a missing value based on the length of the time gap between the two. Although this decay mechanism effectively captures the intuition that more recent observations carry greater diagnostic value, it overlooks domain-specific discrepancies, where different features follow distinct recording frequencies due to clinical practices. This can be illustrated in the example below.
@@ IV-A EHR-Tailored BRITS Adaptations
Let features $f_{1}$ and $f_{2}$ represent heart rate (HR) and systolic blood pressure (SBP), respectively. As a vital sign indicative of a patient’s state, HR is typically monitored at more frequent intervals compared to SBP. BRITS is presented with an observation $\boldsymbol{x_{t}}$ where both HR and SBP values are missing. The time gap vector $\boldsymbol{\delta_{t}}$ shows $\boldsymbol{\delta^{f_{1}}_{t}}=2$ (indicating that the last HR recording occurred 2 time units ago) and $\boldsymbol{\delta^{f_{2}}_{t}}=7$ (indicating that the last SBP recording occurred 7 time units ago).
According to the principle of temporal decay, BRITS would apply a stronger influence from the more recent HR observation compared to the older SBP observation, despite the fact that SBP is typically recorded less frequently and still carries significant diagnostic importance. Consequently, BRITS would incorrectly assign less weight to the last SBP observation compared to HR, overlooking the domain-specific recording patterns, where the last recording of SBP, despite a longer time gap, remains a critical input for imputation.
Our proposed new decay mechanism prioritises recent observations while accounting for the natural variability in healthcare data collection. In addition to using the observed time gap $\boldsymbol{\delta^{d}_{t}}$, we modify the decay function to incorporate the expected time gap $\boldsymbol{\tau}$ between two recordings of a given feature. The expected time gap $\tau_{d}$ for a feature is computed as the median of the time intervals between successive recordings of feature $d$ in the dataset. This adjustment allows the decay function to adapt to the recording patterns of different clinical features, ensuring that features like SBP, which are recorded less frequently, still carry appropriate weight during imputation.
The new decay factor $\gamma_{t}^{d}$ for a feature $d$ at time $t$ is computed as:
$\gamma_{t}^{d}=\exp(-\max(0,W_{\gamma}(\delta_{t}^{d}-\tau_{d})+b_{\gamma}))$ (14)
Where:
  • $\delta_{t}^{d}$ is the time gap since the last observation of feature $d$,
  • $\tau_{d}$ is the median time gap for feature $d$,
  • $W_{\gamma}$ and $b_{\gamma}$ are learnable parameters.
This formulation ensures that the decay factor peaks when the time gap $\delta_{t}^{d}$ closely matches the expected time gap $\tau_{d}$, and declines as the difference between $\delta_{t}^{d}$ and $\tau_{d}$ increases. Thus, observations that fall within their expected time gap contribute more strongly to the imputation process, while those that lie far outside the expected gap are down-weighted as illustrated in the example below.
@@
For the same features $f_{1}$ and $f_{2}$, examining the dataset reveals that the median time gaps for the two features are $\boldsymbol{\tau_{1}}=2$ for HR and $\boldsymbol{\tau_{2}}=10$ for SBP. These medians reflect the fact that HR is routinely monitored more frequently than SBP in clinical practice. By leveraging these median time gaps, the model can account for the different recording frequencies. Here, since the last observed values for both HR and SBP fall within their respective median time gaps, the model assigns comparable importance to both past recordings when imputing the missing values. This approach ensures that the less frequently measured SBP, with its median gap of 10, is not unfairly penalised, preserving the clinical relevance of both features.
Attention-based Hidden State Initialisation: The hidden state of the recurrent component of BRITS are not generated through the raw input. Instead, they receive incomplete data with missingness indicators for imputation, i.e. $\boldsymbol{\hat{h}}$ replaces $\boldsymbol{h}$ in Eq.(5). However in BRITS, the initial hidden state is initialised to zero as in Eq. 3, which causes the model to rely solely on its internal parameters to estimate initial missing values in Eq. 6, ignoring crucial information from prior observations. This can be particularly problematic in medical data, where early data points are often crucial to understanding patient trends, and failure to incorporate them can lead to inaccurate imputations and the loss of important clinical insights as illustrated by the example below.
Example 3.
Consider a patient whose heart rate (HR) and systolic blood pressure (SBP) are being monitored. Over the last few recorded measurements, the patient’s HR has been steadily increasing, indicating a deteriorating condition. However, there is a monitoring gap leading to missing initial values. With the hidden state initialised to zero, BRITS would not only fail to capture the upward trend in HR by disregarding the critical information from later measurements, but could also propagate erroneous assumptions about stability throughout the time series. As the hidden state learns from the imputed values, it would continuously stack these errors, potentially amplifying the misrepresentation of the patient’s condition as stable when in fact urgent intervention is required.
To solve this issue, our approach makes use of the last observed data point and a decay attention mechanism to generate a conditional distribution for the intial hidden state $\boldsymbol{q}(\boldsymbol{h}_{\text{init}}|\boldsymbol{x}_{\text{last\_obs}},\gamma_{t}^{d}))$ within the model distribution $\boldsymbol{p_{\theta}}(\boldsymbol{x_{t}})$. This strategy is designed to provide a more contextually rich starting point for the model, thus enhancing the effectiveness of subsequent imputations. Instead of applying the decay factor directly to the previous hidden state as done in BRITS (Eqs. (5) and Eq.(10)), we use the decay factor to modulate an attention mechanism to better capture long-range and feature-specific temporal dynamics, discussed below.
First, at each time step $s_{t}$, the last observation $\boldsymbol{x_{\text{last\_obs}}}\in\mathbb{R}^{T\times D_{\text{feature}}}$ and the decay factor $\gamma_{t}^{d}$ are projected and encoded to capture their temporal positioning within the sequence:
@@
$\gamma^{{}^{\prime}}_{t}$ $=\text{PosEncoder}(\text{InputProj}(\gamma_{t}^{d}))$ (16)
Then the transformed input representations are concatenated and fed into a Transformer encoder, which is capable of capturing both long-range dependencies and feature-specific interactions:
$\boldsymbol{C}_{\text{in}}=\text{Concat}(\boldsymbol{x^{{}^{\prime}}_{\text{last\_obs}}},\gamma^{{}^{\prime}}_{t})$ (17)
@@
The output of the Transformer is then passed through a series of 1D convolutional layers to adjust the dimensions and initialise the hidden state:
@@
where Eq.(19) transforms $\boldsymbol{C}_{\text{out}}$ from $\mathbb{R}^{2L\times d_{\text{model}}}$ to $\mathbb{R}^{2L\times d_{\text{hidden}}}$ and produces $\boldsymbol{H_{1}}$, Eq.(20) further scales $\boldsymbol{H_{1}}$ to generate the initialised Hidden State $\boldsymbol{h}_{\text{init}}$. This Transformer-based hidden-state initialisation allows the model to better capture the temporal dynamics and variability in feature-recording patterns, providing a more robust foundation for subsequent imputation steps. The CSAI architecture is shown in Figure 3.
Figure 3: The CSAI architecture, which equips BRITS with self-attention. CSAI begins with an input embedding layer, followed by positional embedding to capture time dependencies. These embeddings are processed through multi-head attention, normalization, and feed-forward layers. The output is used to initialize hidden states for subsequent recurrent layers, accounting for both temporal dynamics and domain-specific variability in recording patterns.
IV-B Non-Uniform Masking Strategy
Our non-uniform masking strategy fundamentally diverges from traditional approaches by leveraging the missingness distribution within the dataset. Our strategy is predicated on the principle outlined in the introduction, whereby the probability of missingness in healthcare data is not uniformly distributed across features. Instead, it varies according to healthcare practices and patient conditions. By incorporating this variability into our masking process, we aim to create a more realistic and representative model of the missingness patterns within the dataset.
Our masking algorithm generates masking probabilities using two factors:
  1. wide
    Missingness Distribution $\boldsymbol{P_{\text{dist}}}(d)$: This reflects the likelihood of masking a feature based on its missingness patterns across similar or neighboring observations.
  2. wiide
    Adjustment Factor $R_{\text{factor}}(d)$: A dynamic value that adjusts the masking probability based on how often the feature $d$ is observed. Frequently observed features are masked less to avoid overfitting, while sparsely observed features are masked more to ensure sufficient model learning from limited data points.
For a given feature $d$, the non-uniform masking probability $\boldsymbol{P_{\text{nu}}}(d)$ is determined as follows: